Grade 12Physical Chemistry

No. of oxalic acid molecules in 100 ml of 0.02 N oxalic acid are -
(a) 6.022 x 1020
(b) 6.022 x 1021
(c_)6.022 x 1022
(d) 6.022 x 1023

vardhan
12 Years agoGrade 12

2 Answers

Manas Satish Bedmutha
12 Years ago
As Oxalic acid is dibasic, n-factor (Basicity) = 2
Now, Normality = No. of moles / n-factor / Vol in litres.
So, 0.02 = n x 2/0.1 i.e. n = 0.001
Thus no. of molecules = Avagadro no / No. of moles = 6.022 x 1020 molecules
Nitesh
7 Years ago
Molarity=0.4
Volume of oxalic acid=100ml
So,Molarity=No. Of moles solutie/volume of solution
0.4=no.of moleof solute/100
So, no. Of moles of solute=0.4×100/1000=0.004moles
No.of molecules of oxalic acid=no of moles of solute× avagdro no.
No. Of molecules of oxalic acid=0.004×6.023×[10]23=24.184×[10]23